# Full exceptional collections on Fano threefolds and groups generated by spherical twists on K3 surfaces

### v3

@@ Abstract.
The second author is a senior researcher at the Research Foundation Flanders (FWO). Both authors are also supported by the ERC grant SCHEMES
@@ 1. Introduction
where $[{\mathbb{H}}/F(\delta)]$ is the orbifold quotient of ${\mathbb{H}}^{0}$ by $F(\delta)$.
$[{\mathbb{H}}^{0}/F(\delta)]$ is an orbifold Riemann surface, so in particular its fundamental group is well understood (see §5.2). This yields a good understanding of the group $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$. A similar approach can be applied to subgroups of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$ that contain $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$ such as (the image of) the group ${\mathcal{G}}$ generated by spherical twists, which was introduced above. It turns out that the image $G$ of ${\mathcal{G}}$ in $F(\delta)$ acts freely on ${\mathbb{H}}^{0}$ (see Lemma 5.8(1)), so that ${\mathbb{H}}^{0}/G$ is a non-compact Riemann surface. Hence ${\mathcal{G}}=\pi_{1}([{\mathbb{H}}^{0}/G])$ is actually a free group! Elaborating on this leads to our next main result.
@@ 3. Generalities
We work over the complex numbers. Unless otherwise specified $X$ is a smooth projective variety. We denote by $\mathop{\text{{coh}}}(X)$ the category of coherent sheaves on $X$ and by ${\mathcal{D}}^{b}(X):={\mathcal{D}}^{b}(\mathop{\text{{coh}}}(X))$ the bounded derived category of $\mathop{\text{{coh}}}(X)$. A Fano variety is a smooth projective variety $X$ whose anticanonical divisor $-K_{X}$ is ample. By a K3 surface $Y$ we mean a complex algebraic surface with trivial canonical bundle $\omega_{Y}\simeq{\mathcal{O}}_{Y}$ and $\operatorname{H}^{1}(Y,{\mathcal{O}}_{Y})=0$. We refer the reader to [22] for a comprehensive account on K3 surfaces.
@@
The braid group $B_{n}$ acts on the set of exceptional collections in ${\mathcal{D}}^{b}(X)$ of length $n$ by mutations. More precisely, the standard generator $\sigma_{i}\in B_{n}$, $11\leq i\leq n-$ 1 acts by
@@ 3.1. Spherical twists
[…]
where the right arrow is 00 because $\operatorname{Hom}(S,S[1-\dim X])=0$. This implies that $S$ is a summand of $\operatorname{Hom}^{\bullet}(S,S^{\prime})\otimes S^{\prime}$. Since both $S$ and $S^{\prime}$ are indecomposable, this yields that $S=S^{\prime}[n]$ for some $n\in{\mathbb{Z}}$. ∎
@@
is an equivalence of categories with quasi-inverse given by $-\otimes_{A}^{\mathbb{L}}{\mathcal{E}}$ (see [9]). Denote ¯$\overline{{\mathcal{E}}}:=\bigoplus_{k=1}^{n}i^{*}E_{k}$. By [2, Theorem 3.0.1] and Lemma 3.12 the composition $T_{1}\dots T_{n}$ is the twist along the spherical functor
$-\otimes_{A}^{\mathbb{L}}\overline{{\mathcal{E}}}:{\mathcal{D}}^{b}(A)\to{\mathcal{D}}^{b}(Y).$
@@
We will also consider collections of spherical objects which occur in the following way. Let $Y$ be a K3 surface of degree $2$. Then $Y$ is realised as a double covering $\pi:Y\to{\mathbb{P}}^{2}$ branched along a sextic curve. If $E\in{\mathcal{D}}^{b}({\mathbb{P}}^{2})$ is exceptional, then ${\mathbb{R}}\pi^{*}(E)\in{\mathcal{D}}^{b}(Y)$ is spherical ([49, Proposition 3.13, Example 3.14b]). One also checks that (3.2) holds with the right-hand side replaced by $-\otimes i^{*}\omega_{X}[1]$.
@@ 3.3. Stability conditions and derived autoequivalences of K3 surfaces
[…]
@@
(3.6) $1\to\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)\cong\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\to\operatorname{Aut}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})\to 1$
@@ 3.4. Fuchsian groups
In fact $\operatorname{PSL}_{2}({\mathbb{R}})$ is exactly the Lie group of orientation-preserving isometries of the hyperbolic plane. Note that ${\mathbb{H}}\subset{\mathbb{C}}={\mathbb{A}}^{1}({\mathbb{C}})\subset{\mathbb{P}}^{1}({\mathbb{C}})$. We let $\overline{{\mathbb{H}}}$ be the closure of ${\mathbb{H}}$ in ${\mathbb{P}}^{1}({\mathbb{C}})$. We have ¯$\overline{{\mathbb{H}}}={\mathbb{H}}\cup{\mathbb{P}}^{1}({\mathbb{R}})={\mathbb{H}}\cup{\mathbb{R}}\cup\{\infty\}$. The action of the group $\operatorname{PSL}_{2}({\mathbb{R}})$ extends to $\overline{{\mathbb{H}}}$.
@@
Let $e\mid n$ satisfy $e>0$. An Atkin–Lehner element [^4] Traditionally an Atkin-Lehner element is normalised to have determinant $1$ which involves dividing by $\sqrt{e}$. We do not do this here since we consider $\Gamma^{+}_{0}(n)$ as a subgroup of $\operatorname{PGL}^{+}_{2}({\mathbb{Q}})$. associated to $e$ is a matrix with determinant $e$ of the form
@@
The group $\Gamma_{0}^{+}(n)$ contains $\Gamma_{0}(n)$, and for square free $n$, it is its normaliser. We have $\Gamma^{+}_{0}(n)/\Gamma_{0}(n)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{c}$ where $c$ the number of distinct prime factors in $n$. The groups $\Gamma_{0}^{+}(n)$ and the corresponding modular curves $X_{0}^{+}(n):={\mathbb{H}}/\Gamma_{0}^{+}(n)$ have been studied extensively. For instance, Ogg [41] observed that the primes $p$ for which $X_{0}^{+}(p)$ has genus 00 are precisely the prime factors of the order of the Monster group.
@@ 4.1. Introduction
×${\mathbb{H}}\times\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong{\mathcal{P}}^{+}(Y)$
In particular we may transfer the standard hyperbolic metric on ${\mathbb{H}}$ to ${\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})$. In §4.2,4.3 we give an intrinsic definition for this hyperbolic metric by showing that ${\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})$ may be identified with the hyperboloid model for the hyperbolic plane. We also give a convenient model of ${\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})$ as the open subset ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ of ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})$ consisting of negative vectors for the Mukai quadratic form.
This identification is such that $\overline{\{\rho^{\perp}\mid\rho\in\Delta^{+}(Y)\}}\subset{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})$ is identified with $\{\bar{\rho}\mid\rho\in\Delta^{+}(Y)\}\subset{\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ and the latter is again identified with a discrete set of points in ${\mathbb{H}}$ via the formula (4.8) below. Hence if we use $(-)_{0}$ or $(-)^{0}$ to denote the complement of root data then we have the identifications
@@ 4.2. The hyperboloid model
Let $W$ be a real vector space of dimension ${n+1}$ equipped with a quadratic form $q$ of signature (++,)limit-from($\cdots$+,-)( + ⋯ + , - ). Let $\langle$-,$\rangle$ - , - ⟩ be the corresponding pairing. In particular, $\langle x,x\rangle=2q(x)$. Put
@@ 4.3. Application to K3 surfaces
$\operatorname{Gr}(2,{\mathcal{N}}(Y)_{\mathbb{R}})_{>0}\cong{\mathcal{P}}(Y)/\operatorname{GL}_{2}({\mathbb{R}})\cong{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}}).$
@@
[…]
@@
As explained in §4.2 the composition (4.4) sends $z$ to $w$ where $w\in{\mathbb{R}}^{3}$ is orthogonal to $\operatorname{Re}e(z),\operatorname{Im}e(z)$ and which satisfies $q(w)=-1$, $r>0$. Using (4.2) one checks that one may take $w$ as in (4.5). ∎
@@
[…]
@@ 4.4. Transfer of the $\operatorname{PSL}_{2}({\mathbb{R}})$ action
the two $+$’s have different meanings. In fact (4.10) kills $-1$ (the image of $[1]$) and its image is $\operatorname{SO}^{+}(q,{\mathbb{R}})$. More precisely (4.10) induces an isomorphism
@@
When $\alpha\in{\mathbb{Z}}$ the isometry on ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ induced by $-\otimes{\mathcal{O}}_{Y}(\alpha)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ is given by the formula.
@@ 5.3. Groups generated by point reflections in the hyperbolic plane
[…]
@@ 5.4. Application
      $x$ is a root point of order $>2$.
@@ 6.1. Introduction
The group $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$ acts properly discontinuously (cfr. §3.4) on ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}\cong{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})$ (cfr. §4) and hence also on ${\mathcal{P}}^{+}(Y)$. In particular the quotients
$[{\mathcal{P}}^{+}(Y)/\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})]\text{ and }[{\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}/\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})]$
@@
where $\widetilde{\operatorname{GL}}_{2}^{+}({\mathbb{R}})$ is the universal cover of $\operatorname{GL}_{2}^{+}({\mathbb{R}})$. Since $\widetilde{\operatorname{GL}}^{+}_{2}({\mathbb{R}})$ and $\operatorname{Stab}^{\dagger}(Y)$ are contractible, the same is true for $\operatorname{Stab}^{n}(Y)$.
The element $[1]\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ acts solely by shifting the phases and is contained in $\widetilde{\operatorname{GL}}^{+}_{2}({\mathbb{R}})$ so it acts trivially on $\operatorname{Stab}^{n}(Y)$. Similarly if $d=2$ then we have shown in Lemma B.2 that $\iota$ as introduced in Proposition B.1 acts trivially on $\operatorname{Stab}^{\dagger}(Y)$.
@@
[…]
@@ 6.2. Main result
To simplify the discussion we will identify ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$ and ${\mathbb{H}}$. E.g. when we speak about “roots” in ${\mathbb{H}}$ we mean points of the form $p(\bar{\rho})$ where $\rho\in\Delta^{+}(Y)$. As usual ${\mathbb{H}}^{0}$ is the complement of ${\mathcal{S}}:=\{p(\bar{\rho})\mid\rho\in\Delta^{+}(Y)\}$. Recall that by Lemma 5.8(1), the points of ${\mathcal{S}}$ are order two ramification points for ${\mathbb{H}}\rightarrow{\mathbb{H}}/G$ and furthermore $G$ acts freely on ${\mathbb{H}}^{0}$ so $[{\mathbb{H}}^{0}/G]={\mathbb{H}}^{0}/G$. By Lemma it is sufficient to prove the theorem with ${\mathcal{G}}$ replaced by $\bar{{\mathcal{G}}}$, the image of ${\mathcal{G}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. By Theorem 3.21 $\bar{{\mathcal{G}}}$ contains $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$. By Lemma 6.3 the exact sequence
@@
[…]
@@ 7. Examples
By Theorem 6.5 to find a set of free generators of ${\mathcal{G}}$ in each of the cases it is sufficient to find a set of orbit representatives for the fixed points in ${\mathbb{H}}$ of elements of $G$ contained in an appropriate fundamental domain, where $G\subset\operatorname{Aut}^{+}\operatorname{H}^{*}(Y)$ is the group generated by spherical reflections. The fundamental domains in this section are precisely the “Ford domains” which were defined in the proof of Lemma 3.28. We obtain such a fundamental domain explicitly from a fundamental domain for the corresponding group $\Gamma_{0}^{+}(\delta)$ using Lemma 3.29. In turn, since in all of our cases $\Gamma_{0}^{+}(\delta)=F(\delta)$ is the Fricke group, constructing a fundamental domain for $\Gamma_{0}^{+}(\delta)$ amounts to finding a fundamental domain for $\Gamma_{0}(\delta)$ invariant under the Fricke involution and cutting it along the half-circle with radius $1/\sqrt{\delta}$ centred at 00. These fundamental domains for $\Gamma_{0}^{+}(\delta)$ can be also found in e.g. [5], [51].

### v4

@@ Abstract.
The second author is a senior researcher at the Research Foundation Flanders (FWO). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 885203).
@@ 1. Introduction
where ${\mathbb{H}}/F(\delta)]$ is the orbifold quotient of ${\mathbb{H}}^{0}$ by $F(\delta)$.
${\mathbb{H}}^{0}/F(\delta)]$ is an orbifold Riemann surface, so in particular its fundamental group is well understood (see §5.2). This yields a good understanding of the group $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$. A similar approach can be applied to subgroups of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$ that contain $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$ such as (the image of) the group $\mathcal{G}$} generated by spherical twists, which was introduced above. It turns out that the image $G$ of $\mathcal{G}$} in $F(\delta)$ acts freely on ${\mathbb{H}}^{0}$ (see Lemma 5.8(1)), so that ${\mathbb{H}}^{0}/G$ is a non-compact Riemann surface. Hence ${\mathcal{G}}=\pi_{1}([{\mathbb{H}}^{0}/G])$ is actually a free group! Elaborating on this leads to our next main result.
@@ 3. Generalities
We work over the complex numbers. Unless otherwise specified $X$ is a smooth projective variety. We denote by $\mathop{\text{{coh}}}(X)$ the category of coherent sheaves on $X$ and by ${\mathcal{D}}^{b}(X):={\mathcal{D}}^{b}(\mathop{\text{{coh}}}(X))$ the bounded derived category of $\mathop{\text{{coh}}}(X)$. A Fano variety is a smooth projective variety $X$ whose anticanonical divisor $K_{X}$ is ample. By a K3 surface $Y$ we mean a complex algebraic surface with trivial canonical bundle $\omega_{Y}\simeq{\mathcal{O}}_{Y}$ and $\operatorname{H}^{1}(Y,{\mathcal{O}}_{Y})=0$. We refer the reader to [22] for a comprehensive account on K3 surfaces.
@@
The braid group $B_{n}$ acts on the set of exceptional collections in ${\mathcal{D}}^{b}(X)$ of length $n$ by mutations. More precisely, the standard generator $\sigma_{i}\in B_{n}$, $1\leq i\leq n-1$ acts by
@@ 3.1. Spherical twists
$\operatorname{Hom}^{\bullet}(S,S^{\prime})\otimes S^{\prime}\xrightarrow{}S\xrightarrow{0}S[1-\dim X],$
where the right arrow is 0 because $\operatorname{Hom}(S,S[1-\dim X])=0$. This implies that $S$ is a summand of $\operatorname{Hom}^{\bullet}(S,S^{\prime})\otimes S^{\prime}$. Since both $S$ and $S^{\prime}$ are indecomposable, this yields that $S=S^{\prime}[n]$ for some $n\in{\mathbb{Z}}$. ∎
@@
is an equivalence of categories with quasi-inverse given by $\otimes_{A}^{\mathbb{L}}{\mathcal{E}}$ (see [9]). Denote $\overline{{\mathcal{E}}}:=\bigoplus_{k=1}^{n}i^{*}E_{k}$. By [2, Theorem 3.0.1] and Lemma 3.12 the composition $T_{1}\dots T_{n}$ is the twist along the spherical functor
$\otimes_{A}^{\mathbb{L}}\overline{{\mathcal{E}}}:{\mathcal{D}}^{b}(A)\to{\mathcal{D}}^{b}(Y).$
@@
We will also consider collections of spherical objects which occur in the following way. Let $Y$ be a K3 surface of degree $2$. Then $Y$ is realised as a double covering $\pi:Y\to{\mathbb{P}}^{2}$ branched along a sextic curve. If $E\in{\mathcal{D}}^{b}({\mathbb{P}}^{2})$ is exceptional, then ${\mathbb{R}}\pi^{*}(E)\in{\mathcal{D}}^{b}(Y)$ is spherical ([49, Proposition 3.13, Example 3.14b]). One also checks that (3.2) holds with the right-hand side replaced by $\otimes i^{*}\omega_{X}[1]$.
@@ 3.3. Stability conditions and derived autoequivalences of K3 surfaces
(3.5) ${\mathcal{P}}(Y):=\{v\in{\mathcal{N}}(Y)_{{\mathbb{C}}}\ |\ \operatorname{Re}(v),\operatorname{Im}(v)\text{ span a positive 2-plane}\}$
@@
(3.6) $1\to\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)\cong\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\to\operatorname{Aut}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})\to 1.$
@@ 3.4. Fuchsian groups
[…]
@@
[…]
@@
The group $\Gamma_{0}^{+}(n)$ contains $\Gamma_{0}(n)$, and for square free $n$, it is its normaliser. We have $\Gamma^{+}_{0}(n)/\Gamma_{0}(n)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{c}$ where $c$ the number of distinct prime factors in $n$. The groups $\Gamma_{0}^{+}(n)$ and the corresponding modular curves $X_{0}^{+}(n):={\mathbb{H}}/\Gamma_{0}^{+}(n)$ have been studied extensively. For instance, Ogg [41] observed that the primes $p$ for which $X_{0}^{+}(p)$ has genus 0 are precisely the prime factors of the order of the Monster group.
@@ 4.1. Introduction
${\mathbb{H}}\times\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong{\mathcal{P}}^{+}(Y)$
[…]
@@ 4.2. The hyperboloid model
Let $W$ be a real vector space of dimension n+1{n+1} equipped with a quadratic form $q$ of signature $(+\cdots+,-)$. Let $\langle-,$-$\rangle$ be the corresponding pairing. In particular, $\langle x,x\rangle=2q(x)$. Put
@@ 4.3. Application to K3 surfaces
[…]
@@
(4.4) ${\mathbb{H}}\xrightarrow{e}{\mathcal{P}}^{+}(Y)\rightarrow{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong\operatorname{Gr}(2,{\mathcal{N}}(Y)_{\mathbb{R}})_{>0}\cong{\mathcal{N}}(Y)_{{\mathbb{R}},-1}^{+}$
@@
As explained in §4.2 the composition (4.4) sends $z$ to $w$ where $w\in{\mathbb{R}}^{3}$ is orthogonal to $\operatorname{Re}e(z),\operatorname{Im}e(z)$ and which satisfies q(w)=1q(w)=-1, $r>0$. Using (4.2) one checks that one may take $w$ as in (4.5). ∎
@@
${\mathbb{H}}\xrightarrow{e}{\mathcal{P}}^{+}(Y)\rightarrow{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong\operatorname{Gr}(2,{\mathcal{N}}(Y)_{\mathbb{R}})_{>0}\cong{\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$
@@ 4.4. Transfer of the $\operatorname{PSL}_{2}({\mathbb{R}})$ action
the two ++’s have different meanings. In fact (4.10) kills 1-1 (the image of $[1]$) and its image is $\operatorname{SO}^{+}(q,{\mathbb{R}})$. More precisely (4.10) induces an isomorphism
@@
When $\alpha\in{\mathbb{Z}}$ the isometry on ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ induced by $\otimes{\mathcal{O}}_{Y}(\alpha)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ is given by the formula.
@@ 5.3. Groups generated by point reflections in the hyperbolic plane
(5.4) $1\rightarrow\pi_{1}({\mathbb{H}}-{\mathcal{S}},y)\xrightarrow{\varphi}\pi_{1}(V^{\circ},x)\rightarrow G\rightarrow 1$
@@ 5.4. Application
      $x$ is a root point of order >2>2.
@@ 6.1. Introduction
[…]
@@
[…]
@@
$\operatorname{Stab}^{\dagger}(Y)$${\mathcal{P}}^{+}_{0}(Y)$$\operatorname{Stab}^{n}(Y)$${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}-\overline{\Delta^{+}(Y)}$$\scriptstyle{\widetilde{\operatorname{GL}}^{+}_{2}({\mathbb{R}})}$$\scriptstyle{\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)}$$\scriptstyle{\operatorname{GL}_{2}^{+}({\mathbb{R}})}$$\scriptstyle{\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]}$
@@ 6.2. Main result
To simplify the discussion we will identify ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$ and $\mathbb{H}$}. E.g. when we speak about “roots” in $\mathbb{H}$} we mean points of the form $p(\bar{\rho})$ where $\rho\in\Delta^{+}(Y)$. As usual ${\mathbb{H}}^{0}$ is the complement of ${\mathcal{S}}:=\{p(\bar{\rho})\mid\rho\in\Delta^{+}(Y)\}$. Recall that by Lemma 5.8(1), the points of $\mathcal{S}$} are order two ramification points for ${\mathbb{H}}\rightarrow{\mathbb{H}}/G$ and furthermore $G$ acts freely on ${\mathbb{H}}^{0}$ so ${\mathbb{H}}^{0}/G]={\mathbb{H}}^{0}/G$. By Lemma it is sufficient to prove the theorem with $\mathcal{G}$} replaced by $\bar{{\mathcal{G}}}$, the image of ${\mathcal{G}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. By Theorem 3.21 $\bar{{\mathcal{G}}}$ contains $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$. By Lemma 6.3 the exact sequence
@@
(6.4) $1\rightarrow\pi_{1}({\mathbb{H}}^{0},y)\xrightarrow{\varphi}\pi_{1}({\mathbb{H}}^{0}/G,x)\rightarrow G\rightarrow 1$
@@ 7. Examples
By Theorem 6.5 to find a set of free generators of $\mathcal{G}$} in each of the cases it is sufficient to find a set of orbit representatives for the fixed points in $\mathbb{H}$} of elements of $G$ contained in an appropriate fundamental domain, where $G\subset\operatorname{Aut}^{+}\operatorname{H}^{*}(Y)$ is the group generated by spherical reflections. The fundamental domains in this section are precisely the “Ford domains” which were defined in the proof of Lemma 3.28. We obtain such a fundamental domain explicitly from a fundamental domain for the corresponding group $\Gamma_{0}^{+}(\delta)$ using Lemma 3.29. In turn, since in all of our cases $\Gamma_{0}^{+}(\delta)=F(\delta)$ is the Fricke group, constructing a fundamental domain for $\Gamma_{0}^{+}(\delta)$ amounts to finding a fundamental domain for $\Gamma_{0}(\delta)$ invariant under the Fricke involution and cutting it along the half-circle with radius $1/\sqrt{\delta}$ centred at 0. These fundamental domains for $\Gamma_{0}^{+}(\delta)$ can be also found in e.g. [5], [51].